# FB-Bench: A Fine-Grained Multi-Task Benchmark for Evaluating LLMs’ Responsiveness to Human Feedback

### v1

@@ Abstract
Human feedback is crucial in the interactions between humans and Large Language Models (LLMs). However, existing research primarily focuses on benchmarking LLMs in single-turn dialogues. Even in benchmarks designed for multi-turn dialogues, the user inputs are often independent, neglecting the nuanced and complex nature of human feedback within real-world usage scenarios. To fill this research gap, we introduce FB-Bench, a fine-grained, multi-task benchmark designed to evaluate LLMs’ responsiveness to human feedback in real-world usage scenarios. Drawing from the two main interaction scenarios, FB-Bench comprises 734 meticulously curated samples, encompassing eight task types, five deficiency types of response, and nine feedback types. We extensively evaluate a broad array of popular LLMs, revealing significant variations in their performance across different interaction scenarios. Further analysis indicates that task, human feedback, and deficiencies of previous responses can also significantly impact LLMs’ responsiveness. Our findings underscore both the strengths and limitations of current models, providing valuable insights and directions for future research. Both the toolkits and the dataset of FB-Bench are available at https://github.com/PKU-Baichuan-MLSystemLab/FB-Bench.
@@ 1 Introduction
Equipped with advanced intelligence and formidable processing capabilities, large language models (LLMs) have demonstrated substantial potential extensive potential in seamless interaction with human users and in assimilating real-time human feedback during inference processes [1]. This human-LLM synergy can be mutually beneficial, breaking through the limitations inherent to each side [2, 3] and has been applied in many domains [4, 5, 6, 7].
As a main component of human-LLM synergy, human feedback acts as a response to prior model outputs, serving as a guiding force that directs LLMs towards the desired outcomes [1]. In practical applications, LLMs often need to iteratively adjust their responses based on user feedback in multi-turn dialogues to fulfill user needs. Effective feedback can enhance the quality of responses, while ineffective feedback may mislead LLMs. A robust LLM should leverage appropriate feedback and remain undisturbed by inappropriate feedback. However, evaluating the responsiveness of LLMs to human feedback within multi-turn dialogues presents a significant challenge, as these models exhibit divergent behaviors compared to single-turn dialogues. As illustrated in Figure 1, LLMs that perform well in single-turn interactions may struggle to incorporate user feedback effectively. Conversely, models that may not excel in single-turn scenarios could excel in correcting their previous errors by skillfully leveraging user feedback to improve responses.
@@
Despite the advancement in evaluating LLMs, there remains a lack of a systematic benchmark for evaluating their responsiveness to various human feedback under real-world usage scenarios. Most existing benchmarks [8, 9, 10, 11] assess LLMs within single-turn scenarios, which is naturally incapable of evaluating LLMs’ responsiveness to human feedback. Although there are benchmarks designed for multi-turn dialugues [12, 13, 14], the user inputs are often independent in a dialogue, lacking feedback towards to the previous LLM output. While several benchmarks incorporating feedback exist [15, 16, 17], they typically assess LLMs on a singular task or within a specific domain, and the feedback is often not generated by real humans, failing to capture the complexity and diversity of human-LLM interaction.
In this work, we introduce FB-Bench, a fine-grained multi-task benchmark designed to evaluate LLMs’ responsiveness to various human feedback within real-world usage scenarios. Drawing from the interaction scenarios of error correction and response maintenance, FB-Bench organizes a three-tier hierarchical taxonomy that encapsulates the fundamental elements of human-LLM interaction: user queries, model responses, and user feedback, as illustrated in Figure 2. It includes eight popular task types, five types of deficiencies in previous model outputs, and nine types of user feedback.
After meticulous curation, we collect 734 samples in FB-Bench, each consisting of a task-oriented user query, a preset model response, human feedback, a human-curated reference follow-up response, and a weighted checklist for evaluation. To precisely assess the performance of LLMs in a detailed manner, we employ GPT- 4o to act as a judge, scoring the model-generated follow-up responses based on the weighted checklist and a human-curated reference follow-up response.
We conducted extensive experiments across a broad spectrum of popular LLMs. The results indicate significant variations in performance between error correction and response maintenance scenarios. We further analyze the impact of interaction scenarios, tasks, feedback, and previous responses’ deficiencies on LLMs’ responsiveness. These analyses reveals that
  • Most LLMs exhibit superior performance in error correction compared to response maintenance.
  • Advanced LLMs show similar performance on each task in error correction. In contrast, their performance varies widely in response maintenance.
  • Directly pointing out errors significantly aids LLMs in enhancing the quality of responses, whereas providing feedback with fabricated credentials or expertise often misleads them.
  • Advanced LLMs outperform less capable LLMs in addressing all categories of deficiencies, especially in logic and following instructions.
@@
  • New benchmark. We introduce FB-Bench, the first systematic benchmark for comprehensively evaluating LLMs’ responsiveness to human feedback across a spectrum of real-world, multi-task scenarios.
  • More fine-grained evaluation. We develop a framework that employs a weighted, sample-specific checklist and a human-curated follow-up response to facilitate a fine-grained evaluation of each sample.
  • New findings. We perform a comprehensive evaluation of 31 different LLMs using FB-Bench, uncovering a significant performance discrepancy between error correction and response maintenance. We further analyze the factors that may impact the responsiveness of LLMs and provide valuable insights and directions for future research.
@@ 2 FB-Bench
Figure 2: Overview of FB-Bench. (1)Data Curation: A human-LLM synergy pipeline for mining the target data from real-world usage scenarios and improving their quality and diversity. (2)Three-tier Hierarchical Taxonomy: Comprising 8 popular task types, 2 deficiency types and 9 feedback types, derived from two interaction scenarios. (3)Auto-Evaluation: An LLM-as-a-Judge framework to automatically evaluate LLM’s response with a weighted checklist and a reference follow-up response.
@@ 2.1 Interaction scenario
Response Maintenance: Alternatively, when a user’s query receives an objectively correct or satisfactory response from the model, they might still engage in feedback. This could be to either reaffirm or challenge the provided answer, aiming to verify the correctness and reliability of the information. The expectation is that the model will sustain its initial response upon receiving this feedback.
@@ 2.2.1 Query task
From the perspective of user queries, the diversity of interactions primarily stems from the task type associated with each query. Therefore, we select eight popular tasks to encompass most real-world usage scenarios. To enhance the diversity of queries further, we categorize the eight tasks into twenty-four subtasks, as detailed in Appendix A.1.1.
@@ 2.3 Evaluation Protocol
Inspired by DRFR [18], we evaluate the quality of follow-up responses by decomposing the evaluation criteria into a series of criteria that constitute a checklist. Considering the efficiency and capabilities of LLMs, we adopt the LLM-as-a-Judge framework to evaluate the quality of response as previous works [12, 10]. Specifically, we employ GPT-4o to act as a judge, scoring the model-generated follow-up responses based on the checklist and a human-curated reference follow-up response.
@@
In response maintenance, since the model has already provided the correct answer in the previous round, it will get credits if it maintains its stance and is not swayed by the user’s instructions. That’s to say, meeting any criterion in the checklist yields a score of 1.
@@ 2.4 Dataset Curation
Each sample in FB-Bench mainly contains a task-oriented user query, a preset model response, human feedback, a human-curated reference follow-up response, and a weighted checklist for evaluation.
The example can be found in Appendix A.1.3. The detailed construction pipeline is described as follows.
@@
To ensure the diversity and authenticity of user queries, we mine relevant data from two primary sources: an online chat service and human preference data, both derived from real-world usage scenarios. For error correction data, we employ heuristic rules to identify target data within the online chat service and select the response with the lowest score from human preference data. For response maintenance data, we adopt an opposite strategy to filter the target data from the two data sources. After gathering the above data, we perform deduplication and anonymization, and categorize them into predefined tasks and subtasks using an in-house model to construct high task diversity data.
@@
Although mined data exhibit high task diversity, the feedback from most users is usually simple and homogenous. To improve the quality and variety of user feedback and to supply essential elements for further analysis, we invite annotators to label data with finer granularity. Considering the excellent performance of LLMs in aiding humans to generate comprehensive critiques and reduce hallucination rates [3], we have annotators collaborate with GPT-4 to enhance the quality and efficiency of the annotation process. Initially, we utilize GPT-4 to ascertain the cause of dissatisfaction when a model’s response does not meet the user’s expectations and then simulate a user providing detailed feedback. Subsequently, GPT-4 is tasked with generating a reference follow-up response and a weighted checklist to facilitate the evaluation. Finally, the annotators act as the reviewers to refine all pre-annotated elements of each sample.
@@
To enhance distinguishment in scores among LLMs, we utilize three models: Mistral-7B-Instruct-v0.3 [19], Phi-3-mini-4k-instruct [20], and Yi-1.5-9B-Chat [21] as difficulty filters in our dataset curation pipeline. Specifically, we benchmark these models using this dataset, analyze their responses, and score them by GPT-4o. Finally, we discard samples for which all three models achieved full scores.
@@ 2.5 Dataset Statistics
After meticulous curation, we collect 734 high-quality, diverse, and complex samples. The distributions of tasks, deficiencies in previous model responses within error correction scenarios, and user feedback within both error correction and response maintenance scenarios are all shown in Figure 3. More detailed statistics can be found in Appendix A.1.2
Figure 3: FB-Bench Statistics.
@@ 3.1 Experimental Setup
Models. Given the considerations of performance, size and the popularity of LLMs, we systematically evaluate a wide array of LLMs, including GPT family, Claude-3.5, Qwen-2.5 family, ERNIE-4, Moonshot, Yi, Gemma-2, Mistral, InternLM-2.5, DeepSeek, GLM-4, Phi-3 and LlaMa-3.1 family [22, 23, 24, 25, 19, 26, 27, 28, 20, 29].
Response generation. We employ the official settings and chat template in HuggingFace model card for open-source LLMs. Proprietary models are assessed via their official API endpoints, using the latest versions as of September 25, 2024. Considering the varied requirements for diversity and creativity across tasks, we set different temperatures for different tasks. More details can be found in Appendix A.2.1.
Evaluation. We utilize gpt-4o-2024-08-06 to evaluate each generated follow-up response using the weighted checklist and the corresponding reference follow-up response. The evaluation prompt and cases can be found in Appendix A.2.2. To enhance the determinism of the judgment, we set the temperature to 0 and the output length to 4096.
@@ 3.2 Main Results
  • The ranking of LLMs exhibits significant variation between error correction and response maintenance, indicating an unrelated relationship between these two capabilities. For error correction, the ranking of LLMs aligns with people’s perceptions, where closed-source LLMs significantly outperform open-source LLMs. The top three LLMs are closed-source, and Qwen2.5-72B-Instruct achieves the highest score among open-source LLMs but still lags significantly behind the leading closed-source LLM. Conversely, in response maintenance scenarios, the top four LLMs comprise two open-source LLMs: Qwen2.5-72B-Instruct and internlm2_5-20b-chat, with the former achieving the highest performance among all LLMs.
  • Some LLMs perform well in error correction yet struggle in response maintenance, indicating they are more susceptible to unreasonable user feedback. LLMs such as Deepseek-V2.5, claude-3-5-sonnet-20240620 and gpt-4o-mini-2024-07-18 exemplify this trend. Specifically, claude-3-5-sonnet-20240620 attains the highest scores in error correction but demonstrates relatively weak response maintenance capability, resulting in its lower ranking.
Figure 4: The subset evaluation results in FB-Bench between error correction and response maintenance scenarios. Overall denotes the mean of error correction score and response maintenance score.
@@ 3.3 Analysis
Thanks to our comprehensive taxonomy, we can delve into several critical factors that significantly influence the performance of LLMs on FB-Bench, including interaction scenario types, task types, feedback types and deficiency types.
Most LLMs exhibit superior performance in error correction compared to response maintenance.
The performance discrepancy between these two scenarios is illustrated in Figure 5. Generally, LLMs excel in error correction but struggle to maintain consistency in their responses. This disparity is more pronounced among LLMs developed outside of China, even for leading LLMs such as gpt-4o-2024-05-13 and claude-3.5-sonnet-20240620. This suggests that the majority of LLMs lack a robust capacity to differentiate between valid and misleading instructions. This deficiency could stem from the fact that optimizing an LLM’s instruction-following ability is relatively straightforward; however, overly adherent instruction-following can cause the LLM to distort reality.
Figure 5: Performance discrepancy between error correction and response maintenance.
@@
Advanced LLMs show similar performance on each task in error correction. In contrast, their performance varies widely in response maintenance.
@@
We present the performance scores of several advanced LLMs across different tasks in Figure 6(a) and Figure 6(b). In error correction scenarios, the scores of different LLMs on each task are relatively close, and they exhibit notably poorer performance on mathematics and reasoning tasks, with scores hovering around 60. Conversely, in response maintenance scenarios, the score discrepancies among different LLMs on each task are more pronounced, particularly in mathematics and reasoning tasks. Remarkably, claude-3-5-sonnet-20240620’s ability in response maintenance is significantly weaker compared to other LLMs, especially in reasoning tasks, despite achieving the highest score in the error correction dimension.
Figure 6: (a) The performance of 4 top-tier LLMs across 8 popular tasks within error correction scenarios. (b) The performance of 4 top-tier LLMs across 8 popular tasks within response maintenance scenarios. (c) The performance of 4 vastly different LLMs across five types of discrepancies in previous responses within error correction scenarios.
@@
Pointing out errors directly significantly helps LLMs enhance the quality of responses, while challenging LLMs with fabricated credentials or expertise often mislead them.
@@
We present the performance of several advanced LLMs under various types of human feedback in Figure 7. In error correction scenarios, all LLMs achieve scores higher than 70 when errors are identified by humans. In response maintenance scenarios, all LLMs exhibit poor performance when challenged by humans with fabricated credentials or expertise. Furthermore, it is observable that claude-3-5-sonnet-20240620 significantly outperforms all other LLMs when receiving unreasonable requests, indicating its superior safety capabilities.
Figure 7: Impact of different feedback types.
Advanced LLMs outperform less capable LLMs in addressing all categories of deficiencies, particularly in logic and following instructions.
@@
To deeply investigate the performance disparities in error correction among models, we selected four models that exhibit significant differences in this aspect. Their performance across various deficiency types is depicted in Figure 6(c). It shows that more advanced LLMs outperform less capable ones in all categories of deficiency. The primary challenges identified include correcting logical errors and following user instructions, where smaller models underperform even after receiving human feedback. This underperformance is likely attributable to the limited capabilities of smaller models.

### v2

@@ Abstract
Human feedback is crucial in the interactions between humans and Large Language Models (LLMs). However, existing research primarily focuses on benchmarking LLMs in single-turn dialogues. Even in benchmarks designed for multi-turn dialogues, the user inputs are often independent, neglecting the nuanced and complex nature of human feedback within real-world usage scenarios. To fill this research gap, we introduce FB-Bench, a fine-grained, multi-task benchmark designed to evaluate LLMs’ responsiveness to human feedback under real-world usage scenarios in Chinese. Drawing from the two main interaction scenarios, FB-Bench comprises 591 meticulously curated samples, encompassing eight task types, five deficiency types of response, and nine feedback types. We extensively evaluate a broad array of popular LLMs, revealing significant variations in their performance across different interaction scenarios. Further analysis indicates that task, human feedback, and deficiencies of previous responses can also significantly impact LLMs’ responsiveness. Our findings underscore both the strengths and limitations of current models, providing valuable insights and directions for future research. Code and datasets are available at https://github.com/PKU-Baichuan-MLSystemLab/FB-Bench.
FB-Bench: A Fine-Grained Multi-Task Benchmark for Evaluating LLMs’ Responsiveness to Human Feedback
Youquan Li1, Miao Zheng2, Fan Yang2, Guosheng Dong2, Bin Cui1, Weipeng Chen2, Zenan Zhou2thanks: Corresponding author., Wentao Zhang [^1] footnotemark: 1 1 Peking University 2 Baichuan Inc. youquan.li@stu.pku.edu.cn, zhouzenan@baichuan-inc.com, wentao.zhang@pku.edu.cn
@@ 1 Introduction
Equipped with advanced intelligence and formidable processing capabilities, large language models (LLMs) have demonstrated extensive potential in seamless interaction with human users and in assimilating real-time human feedback during inference processes (Fernandes et al., 2023). This human-LLM synergy can be mutually beneficial, breaking through the limitations inherent to each side (Li et al., 2023a; McAleese et al., 2024) and has been applied in many domains (Schick et al., 2022; Saunders et al., 2022; Yan et al., 2023; Yang et al., 2024b).
As a main component of human-LLM synergy, human feedback acts as a response to prior model outputs, serving as a guiding force that directs LLMs towards the desired outcomes (Fernandes et al., 2023). In practical applications, LLMs often need to iteratively adjust their responses based on user feedback in multi-turn dialogues to fulfill user needs. Effective feedback can enhance the quality of responses, while ineffective feedback may mislead LLMs. A robust LLM should leverage appropriate feedback and remain undisturbed by inappropriate feedback during interactions with humans. However, evaluating the responsiveness of LLMs to human feedback within multi-turn dialogues presents a significant challenge, as these models exhibit divergent behaviors compared to single-turn dialogues. As illustrated in Figure 1, LLMs that perform well in single-turn interactions may struggle to incorporate user feedback effectively. Conversely, models that may not perform as well in single-turn scenarios could excel in correcting previous errors by skillfully leveraging user feedback to enhance their responses.
@@
Despite the advancement in evaluating LLMs, there remains a lack of a systematic benchmark for evaluating their responsiveness to various human feedback under real-world usage scenarios. Most existing benchmarks (Li et al., 2024; Liu et al., 2023a; Li et al., 2023b; Ni et al., 2024) assess LLMs within single-turn scenarios, which are naturally incapable of evaluating LLMs’ responsiveness to human feedback. Although there are benchmarks designed for multi-turn dialogues (Zheng et al., 2024; Sun et al., 2024; Bai et al., 2024), the user feedback is typically independent of the previous LLM output in a dialogue. While several benchmarks incorporating feedback exist (Wang et al., 2023; Yang et al., 2024a; Liu et al., 2023b), they typically assess LLMs on a singular task or within a specific domain, and the feedback is often not generated by real humans, failing to capture the nuance and diversity of human-LLM interaction.
In this work, we introduce FB-Bench, a fine-grained multi-task benchmark designed to evaluate LLMs’ responsiveness to various human feedback under real-world usage scenarios in Chinese. Drawing from the interaction scenarios of error correction and response maintenance, FB-Bench organizes a three-tier hierarchical taxonomy that encapsulates the fundamental elements of human-LLM interaction: user queries, model responses, and user feedback, as illustrated in Figure 2. It includes eight popular task types, five types of deficiencies in previous model outputs, and nine types of user feedback.
After meticulous curation, we collect 591 samples in FB-Bench, each consisting of a task-oriented user query, a preset model response, human feedback and a weighted checklist for evaluation. To precisely assess the performance of LLMs in a detailed manner, we employ GPT- 4o to act as a judge, scoring the model-generated follow-up responses based on the human-curated weighted checklist. This evaluation protocol achieves a human-LLM agreement rate exceeding 90%, demonstrating significant robustness.
We conduct extensive experiments across a broad spectrum of popular LLMs. The results indicate that the performance gap between open-source and closed-source LLMs is narrowing. Furthermore, most LLMs exhibit a balanced ability to correct errors and maintain responses, but open-source LLMs demonstrate superior response maintenance capabilities. We further analyze the impact of tasks, feedback, and previous responses’ deficiencies on LLMs’ responsiveness. These analyses reveals that
  • Leading LLMs show similar performance on each task in error correction. In contrast, their performance varies significantly in response maintenance.
  • Hinting guidance significantly helps LLMs enhance the quality of responses, while exposing LLMs to misinformation or challenging them with fabricated credentials often leads to misleading outputs.
  • Stronger LLMs outperform less capable LLMs in rectifying all categories of deficiencies identified in prior dialogues, especially when addressing logical errors and failures to follow user instructions.
@@
  • New benchmark. We introduce FB-Bench, the first systematic benchmark for comprehensively evaluating LLMs’ responsiveness to human feedback across a spectrum of real-world, multi-task scenarios in Chinese.
  • More fine-grained evaluation. We develop a framework that employs a sample-specific weighted checklist to facilitate a fine-grained evaluation of each sample.
  • New findings. We perform a comprehensive evaluation of 27 different LLMs using FB-Bench, uncovering a significant performance discrepancy between error correction and response maintenance scenarios. We further analyze the factors that may impact the responsiveness of LLMs and provide valuable insights and directions for future research.
@@ 2 FB-Bench
Figure 2: Overview of FB-Bench. (1)Data Curation: A human-LLM synergy pipeline for mining target data from real-world scenarios and improving their quality and diversity. (2)Three-tier Hierarchical Taxonomy: Comprising 8 popular task types, 5 deficiency types and 9 feedback types, derived from two interaction scenarios. (3)Auto-Evaluation: A LLM-as-a-Judge framework to automatically evaluate LLM’s response with a weighted checklist.
@@ 2.1 Interaction scenario
Response Maintenance: Alternatively, when a user’s query receives an objectively correct or satisfactory response from the model, users might still engage in feedback. This could be to either reaffirm or challenge the provided answer, aiming to verify the correctness and reliability of the information. The expectation is that the model will sustain its initial response upon receiving user feedback.
@@ 2.2.1 Query task
From the perspective of user queries, the diversity of interactions primarily stems from the task type associated with each query. Therefore, we select eight popular tasks to encompass most real-world usage scenarios. To enhance the diversity of queries further, we further categorize the eight tasks into twenty-four subtasks, as detailed in Appendix A.1.1.
@@ 2.3 Evaluation Protocol
Inspired by DRFR (Qin et al., 2024), we evaluate the quality of models’ follow-up responses by decomposing the evaluation criteria into a series of criteria that constitute a checklist. Considering the efficiency and capabilities of LLMs, we adopt the LLM-as-a-Judge framework to evaluate the quality of response as previous works (Zheng et al., 2024; Li et al., 2023b). Specifically, we employ GPT-4o to act as a judge, scoring the model-generated follow-up responses based on the human-curated checklist.
@@
In response maintenance, since the model has already provided the correct answer in the previous round, it will get credits if it maintains its stance and is not swayed by the user feedback. That’s to say, meeting any criterion in the checklist yields a score of 1.
@@ 2.4 Dataset Curation
Each sample in FB-Bench mainly contains a task-oriented user query, a preset model response, human feedback and a human-curated weighted checklist for evaluation. The example can be found in Appendix A.1.3. The detailed construction pipeline is described as follows.
@@
To ensure the diversity and authenticity of user queries, we mine relevant data from two primary sources: an online chat service and human preference data. Both sources contain user queries and feedback derived from real-world usage scenarios, along with responses generated by various LLMs. For error correction data, we employ heuristic rules to identify target data within the online chat service and select the response with the lowest score from human preference data. For response maintenance data, we adopt an opposite strategy to filter the target data from the two data sources. The detailed heuristic rules can be found in Appendix  A.2.1. After gathering the above data, we perform deduplication and anonymization, and categorize them into predefined tasks and subtasks using an in-house model to construct high task diversity data.
@@
Although mined data exhibit high task diversity, the feedback from most users is usually simple and homogenous. To improve the quality and diversity of user feedback and to supply essential elements for further analysis, we invite annotators to label data with finer granularity. Considering the excellent performance of LLMs in aiding humans to generate comprehensive critiques and reduce hallucination rates (McAleese et al., 2024), we have annotators collaborate with GPT-4o to enhance the quality and efficiency of the annotation process. Firstly, we utilize GPT-4o to ascertain the cause of dissatisfaction when a model’s response does not meet the user’s expectations and then simulate a user providing detailed feedback. Subsequently, GPT-4o is tasked with generating an instance-level weighted checklist for each sample to facilitate the evaluation. Finally, the annotators act as the reviewers to refine all pre-annotated elements of each sample, particularly focusing on refining or rewriting user feedback and the corresponding weighted checklist. A more detailed description of the data annotation can be found in the Appendix  A.2.2.
Figure 4: The subset evaluation results in FB-Bench between error correction and response maintenance scenarios. Overall denotes the mean of error correction score and response maintenance score. The dashed line represents the diagonal $y=x$.
@@
To enhance distinguishment in scores among LLMs, we utilize three models, including Meta-Llama-3.1-8B-Instruct (Dubey et al., 2024), Phi-3-medium-4k-instruct (Abdin et al., 2024b), and Yi-1.5-9B-Chat (Young et al., 2024) as difficulty filters in our dataset curation pipeline. These models, with their diverse architectures and capabilities, provide a comprehensive assessment of the dataset’s difficulty. Specifically, we benchmark these models using this dataset, analyze their responses, and score them by GPT-4o. Finally, we discard samples for which all three models achieved full scores. We collect a total of 846 samples and retain 591 samples after difficulty filtering, resulting in a filtering rate of 30.14%.
@@ 2.5 Dataset Statistics
After meticulous curation, we collect 591 high-quality, diverse, and complex samples. The distributions of tasks, deficiencies in previous model responses within error correction scenarios, and user feedback within both error correction and response maintenance scenarios are all shown in Figure 3. More detailed statistics can be found in Appendix A.1.2
@@ 3.1 Experimental Setup
Models. Given the considerations of performance, size and the popularity of LLMs, we systematically evaluate a wide array of LLMs, including GPT family, Claude-3.5, Qwen-2.5 family, ERNIE-4, Moonshot, Yi, Gemma-2, Mistral, InternLM, DeepSeek, GLM-4, Phi and LlaMa-3.1 family (Achiam et al., 2023; Team, 2024; moo, ; Team et al., 2024; Jiang et al., 2023; Cai et al., 2024; Liu et al., 2024; GLM et al., 2024; Abdin et al., 2024b, a; Dubey et al., 2024).
Response generation. We employ the official settings and chat template in HuggingFace model card for open-source LLMs. Proprietary models are assessed via their official API endpoints. Considering the varied requirements for diversity and creativity across tasks, we set different temperatures for different tasks. More details can be found in Appendix A.3.1.
Evaluation. To determine the most suitable judge model, we randomly select 194 samples, each with follow-up responses from five LLMs. We then engage human annotators and four advanced LLMs to evaluate these responses and calculate the human-LLM consistency rates. We found that gpt-4o-2024-08-06 achieved the highest consistency rate, at 90.91%. The detailed experimental setup and results can be found in Appendix A.4. Consequently, we choose gpt-4o-2024-08-06 as judge model to evaluate each generated follow-up response based on the corresponding weighted checklist. To enhance the determinism of the judgment, we set the temperature to 0 and the output length to 4096. The evaluation prompt and cases are presented in Appendix A.3.2.
@@ 3.2 Main Results
The performance gap between open-source LLMs and closed-source LLMs is narrowing. DeepSeek-V3 ranks first in overall performance, followed by qwen-max-0919 and gpt-4o-2024-11-20, demonstrating the best error correction ability and third in response maintenance ability. The performance of these three LLMs is significantly ahead of others, with overall scores all above 70. The fourth-ranked LLM, Qwen2.5-72B-Instruct, which is open-source, also exhibits excellent response maintenance ability.
Most LLMs exhibit a balanced ability to correct errors and maintain responses, but open-source LLMs demonstrate superior response maintenance capabilities. Most models are positioned near the diagonal $y=x$, indicating their balanced proficiency in error correction and response maintenance. However, some open-source LLMs are situated above the diagonal $y=x$, suggesting they excel more in maintaining responses than in correcting errors. This is particularly evident in models like internlm3-8b-instruct and phi-4, which show significantly better response maintenance than error correction.
@@ 3.3 Analysis
Thanks to our comprehensive taxonomy, we can delve into several critical factors that significantly influence the performance of LLMs on FB-Bench, including task types, feedback types and deficiency types.
@@
Leading LLMs show similar performance on each task in error correction. In contrast, their performance varies significantly in response maintenance.
@@
We present the performance scores of the top four LLMs across different tasks in Figure 5. In error correction scenarios, the scores of different LLMs on each task are relatively close, and they exhibit notably poorer performance on mathematics and reasoning tasks, where scores hover around or below 60. Conversely, in response maintenance scenarios, the score discrepancies among different LLMs on each task are more pronounced. Specifically, in reasoning tasks, the performances of the four LLMs vary significantly, with gpt-4o-2024-11-20 lagging considerably behind the other three. Additionally, Qwen-2.5-72B-Instruct also falls significantly behind the other three LLMs in text translation tasks.
Figure 5: The performance of top four LLMs across eight popular tasks
@@
Hinting guidance significantly helps LLMs enhance the quality of responses, while exposing LLMs to misinformation or challenging them with fabricated credentials often leads to misleading outputs.
@@
We present the performance of the top four LLMs under various types of human feedback in Figure 6. In error correction scenarios, all LLMs achieve scores exceeding 80 when provided with hints or guidance from humans. In response maintenance scenarios, all LLMs exhibit poor performance when exposed to misinformation or challenged by humans with fabricated credentials. Furthermore, it is evident that qwen-max-0919 significantly outperforms all other LLMs when handling unreasonable requests, indicating its superior safety capabilities.
Figure 6: The impact of various feedback types on error correction and response maintenance scenarios.
Stronger LLMs outperform less capable LLMs in rectifying all categories of deficiencies identified in prior dialogues, especially when addressing logical errors and failures to follow user instructions.
@@
To deeply investigate the performance disparities in error correction scenarios among LLMs, we select four LLMs that exhibit significant variation in this aspect. Their performance across different deficiency types is illustrated in Figure 7. The results indicate that stronger LLMs consistently outperform less capable ones across all deficiency categories. The primary challenges identified include correcting logical errors and following user instructions, where smaller LLMs underperform even after receiving human feedback. This underperformance is likely attributable to the limited capabilities of smaller LLMs.
Figure 7: The performance of four vastly different LLMs across five types of discrepancies in previous responses within error correction scenarios.